20c+10m=495

8c+m=178.50


1
Expert's answer
2020-10-28T17:49:21-0400

20c+10m=495

8c+m=178.50


M = 178.50- 8c


20c + 10(178.50-8c) =495


20c + 1785-80c = 495


20c - 60c = 495- 1785


-40c = -1290


C = 32.25



M = 178.50- 8c


M =178.50- 8(32.25)


M = 178.50- 258

M = -79.5

C = 32.25